# Risk-Aware Distributional Intervention Policies for Language Models

### v1

@@ Abstract.
Language models are prone to occasionally undesirable generations, such as harmful or toxic content, despite their impressive capability to produce texts that appear accurate and coherent. This paper presents a new two-stage approach to detect and mitigate undesirable content generations by rectifying activations. First, we train an ensemble of layerwise classifiers to detect undesirable content using activations by minimizing a smooth surrogate of the risk-aware score. Then, for contents that are detected as undesirable, we propose layerwise distributional intervention policies that perturb the attention heads minimally while guaranteeing probabilistically the effectiveness of the intervention. Benchmarks on several language models and datasets show that our method outperforms baselines in reducing the generation of undesirable output. Our code is available at https://github.com/nguyenngocbaocmt02/OT-Intervention
The authors are with the Chinese University of Hong Kong (nbnguyen, nguyen@se.cuhk.edu.hk), the National University of Singapore (binhnt@nus.edu.sg), and the University of North Carolina at Chapel Hill (duykhuongnguyen277@gmail.com).
1. Introduction
@@
Language models (LMs) have demonstrated remarkable ability to understand and generate human-like documents [35, 6, 42, 43, 20, 11]. However, inspecting their outputs can often reveal undesirable content, such as inaccurate or toxic generated texts [19, 36, 48]. Meanwhile, devising good strategies to control the LMs’ generation process remains a challenge [41]. Numerous methods have been proposed for controllable text generation in language models; see, for example, [51] and [23]. These approaches include model editing and supervised fine-tuning. However, both approaches require altering the model weights using a subset of text samples, which can result in unstable representations for other text instances [16]. In addition, these methods typically require substantial computational resources.
@@
To resolve these issues, one possible alternative for controllable text generation is activation intervention [38, 17, 24], where one alters the model activations responsible for the undesirable output during inference. Previous work highlighted the presence of interpretable directions within the activation space of language models. These directions have been shown to play a causal role during inference. For instance, [7] and [29] suggest that these directions could be manipulated to adjust the model behavior in a controlled manner. This line of work indicates that the internal representations of language models are structured in ways that can be leveraged for fine-grained control over generated text. Taking inspiration from these previous works, activation intervention frameworks argued that the information needed to steer the model to generate a target sentence is already encoded within the model. The hidden information is extracted as latent vectors and then used to guide the generation to have desirable effects. The preliminary success of these activation intervention methods motivates our approach to improve the desirable generation of LMs.
@@
Problem Statement. We consider a language model consisting of $L$ layers, each layer has $H$ head, each head has dimension $d$. For example, for the Llama-2, we have $L=32$, $H=32$ and $d=128$. The training dataset is denoted by $\mathcal{D}=(x_{i},y_{i}^{*})_{i=1,\ldots,N}$, the $i$-th text is denoted by $x_{i}$, and its ground truth label is $y_{i}^{*}\in\{0,1\}$, where the label 1 (positive) represents the undesirable text, and the label 0 (negative) represents the desirable text. Our goal is two-fold: (i) detect an undesirable text, and (ii) modify an undesirable text into a desirable text.
@@
$a_{\ell+1,i}=a^{\mathrm{mid}}_{\ell,i}+\mathrm{FFN}(a^{\mathrm{mid}}_{\ell,i}),\quad a^{\mathrm{mid}}_{\ell,i}=a_{\ell,i}+\sum_{i=1}^{H}Q_{\ell h}\textrm{Att}(P_{\ell h}a_{\ell,i}).$ (1)
Here, $P_{\ell h}\in\mathbb{R}^{d\times dH}$ is the projection matrix that maps the output of each layer to the $d$ dimensional head space, $\mathrm{Att}$ is the attention operator [45], $Q_{\ell h}\in\mathbb{R}^{dH\times d}$ is the pull-back matrix and $\mathrm{FFN}$ is the feed-forward layer. Each $a_{\ell,i}$ is a concatenation of headwise activations $a_{\ell h,i}$ for $h=1,\ldots,H$. Inspired by [24], we aim to perform intervention at some selected $a_{\ell h,i}$, the activations for head $h$ of layer $\ell$, if we detect that the activation is from an undesirable content.
Contributions. We contribute a novel activation intervention method to detect and rectify the undesirable generation of LM. We call our method RADIANT (Risk-Aware Distributional Intervention Policies for Language Models’ Activations). Overall, RADIANT comprises two components:
  1. (1)
@@ 1.1. Related Works
Controllable generation. Controllable text generation methods aim to alter the outputs of large language models in a desired way. One possible approach is model editing [46, 52], which involves modifying the parameters of a model to steer its outputs. For example, [28] involves identifying specific middle-layer feed-forward modules that correspond to factual knowledge and then altering these weights to correct or update the information encoded by the model. Other notable methods include fine-tuning techniques such as Supervised Fine-Tuning (SFT, [32, 14]) and Reinforcement Learning from Human Feedback (RLHF, [30, 13]).
@@
Probing. Probing is a well-established framework to assess the interpretability of neural networks [1, 4]. Probing techniques have been applied to understand the internal representations of transformer architectures in language models such as BERT and GPT. For example, [7] proposed an unsupervised probing method that optimizes consistency between positive and negative samples. [27] computes the mean difference between true and false statements and skews the decision boundary by the inverse of the covariance matrix of the activations.
@@
Activation interventions. Activation intervention at inference time is an emerging technique for controllable generation [44, 24, 37, 50]. Unlike model editing or fine-tuning techniques, inference time intervention does not require altering the model parameters. [24] proposed a headwise intervention method for eliciting truthful generated answers of a language model. They first train linear probes on each head of the language model, then shift the activations with the probe weight direction or mean difference direction.
@@
There is a clear distinction between our method and ITI when choosing the location of the classifiers and, hence, the location of the interventions. The ITI method builds different headwise classifiers scattered at different layers, and it may suffer from distribution shifts: if an activation is intervened, this leads to shifts in the activation values at all subsequent layers in the network. Thus, classifiers trained at subsequent layers can degrade performance, and interventions at subsequent layers can also degrade. On the contrary, we build a layerwise classifier focusing on all heads in the same layer and does not suffer from the distributional shifts of the activations.
@@
The recent paper by [37] is closely related to our work. The authors propose a heuristic intervention rule; then, using empirical estimations of the means and covariances of activations data’s distributions of desirable and undesirable text, they calculate a closed-form optimal transport plan between these two empirical distributions, assuming they are standard normal. However, this framework does not take into account the semantics of sentences. Another recent method, called LoFit (Localized Fine-Tuning on LLM Representations [50]), also identifies a specific subset of attention heads that are crucial for learning a particular task, but then performs fine-tuning on the intervention vectors at those chosen heads to enhance the model’s hidden representations. This results in additional training overhead.
2. Layerwise Risk-aware Probes
@@
$\mathcal{C}_{\ell h}(a_{\ell h})=\begin{cases}1&\text{if }\mathrm{sigmoid}(\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h})\geq 0.5,\\ 0&\text{otherwise,}\end{cases}=\begin{cases}1&\text{if }\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h}\geq 0,\\ 0&\text{if }\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h}<0.\end{cases}$
The training process of $\mathcal{C}_{\ell h}$ must take into account two types of risk: (i) false-negative risk when an undesirable text is not detected, (ii) false-positive risk when a desirable text is classified as undesirable, and is subsequently edited and loses its original semantics. Therefore, a natural candidate for the loss function is a combination of the False Positive Rate (FPR) and the False Negative Rate (FNR). However, neither FPR nor FNR have smooth functions in optimizing variables. We, hence, resort to smooth surrogates of these two metrics that use the predicted probability of the classifier, similarly to [5]. In detail, we use
$\mathrm{FPR}(\theta_{\ell h},\vartheta_{\ell h})$ $=\frac{1}{N_{0}}\sum_{i=1}^{N}\mathrm{sigmoid}(\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h,i})\times(1-y^{*}_{i}),$
$\mathrm{FNR}(\theta_{\ell h},\vartheta_{\ell h})$ $=\frac{1}{N_{1}}\sum_{i=1}^{N}\left(1-\mathrm{sigmoid}(\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h,i})\right)\times y^{*}_{i}.$
@@
where $\tau\in[0,H]$ is a tunable threshold. When $\tau=\lfloor H/2\rfloor$, then $\mathcal{C}_{\ell}$ becomes the majority voting results of the individual (weak) classifiers $\mathcal{C}_{\ell h}$. We optimize the hyperparameter $\tau$ to reduce the False Negative Rate (FNR), with a secondary focus on the False Positive Rate (FPR) in cases of equal FNR rates. The rationality for this choice is that we believe undesirable content being labeled as desirable is more problematic than other instances.
@@
To conclude this step, we can compute the classifier $\mathcal{C}_{\ell}$ for each layer $\ell=1,\ldots,L$ by tuning the parameters $(\alpha)$. The layer whose classifier $\mathcal{C}_{\ell}$ delivers the highest quality (accuracy or any risk-aware metric) will be the optimal layer to construct the probe. This optimal layer, along with the collection of headwise classifiers, is the final output of this step.
(b) FNR across layers for different value of regularization parameter $\alpha$ of the risk-aware loss Eq (2).
Figure 1. Plot of different risk-aware metrics (FNR and FPR) with different values of hyperparameters $\alpha$ across layers of Llama-7B.
3. Headwise Interventions with Probabilistic Guarantees
@@
We propose a distributional intervention to the activations of the samples predicted undesirable by the layerwise classifier. In this section, we will focus on constructing a single headwise intervention, and in the next section, we will combine multiple headwise interventions into a layerwise intervention. A headwise intervention is a map $\Delta_{\ell h}:a_{\ell h}\mapsto\hat{a}_{\ell h}$ that needs to balance multiple criteria: (i) it should be easy to compute and deploy, (ii) it should be effective in converting the undesirable activations to the desirable regions, (iii) it should minimize the magnitude of the intervention to sustain the context of the input. Intuitively, we propose solving an optimization problem with the loss and constraints that fit all the criteria listed. The details are as follows.
@@
To promote (i), we employ a simple linear map $\Delta_{\ell h}(a_{\ell h})=G_{\ell h}a_{\ell h}+g_{\ell h}$ parametrized by a matrix $G_{\ell h}\in\mathbb{R}^{d\times d}$ and a vector $g_{\ell h}\in\mathbb{R}^{d}$. This linear map can also be regarded as a pushforward map that transforms the undesirable-predicted activations to become desirable-predicted activations. Let us now represent the undesirable-predicted activations as a $d$-dimensional random vector $\tilde{a}_{\ell h}$. Its distribution can be estimated using the training data after identifying the subset $\hat{\mathcal{D}}_{\ell h}^{+}$ of training samples that are predicted undesirable by $\mathcal{C}_{\ell h}$, that is, $\hat{\mathcal{D}}_{\ell h}^{+}\triangleq\{i:\mathcal{C}_{\ell h}(a_{\ell h,i})=1\}$. The activations of samples in $\hat{\mathcal{D}}_{\ell h}^{+}$ lead to an empirical distribution ^$\widehat{\mathbb{P}}_{\ell h}$. The linear map $\Delta_{\ell h}$ will pushforward the distribution ^$\widehat{\mathbb{P}}_{\ell h}$ to the new distribution $\mathbb{Q}_{\ell h}=\Delta_{\ell h}\#\widehat{\mathbb{P}}$.
@@
Using the pushforward distribution $\mathbb{Q}_{\ell h}$, we can impose criteria (ii) and (iii) above in an intuitive method. To promote (ii), we require that the activations distributed under $\mathbb{Q}_{\ell h}$ should be classified as desirable by $\mathcal{C}_{\ell h}$ with high probability. Finally, to promote (iii), we require that the distributions $\mathbb{Q}_{\ell h}$ and ^$\widehat{\mathbb{P}}_{\ell h}$ be not too far from each other. Let $\gamma\in(0,0.5)$ be a small tolerance parameter, and let $\varphi$ be a measure of dissimilarity between probability distributions, we propose to find $\Delta_{\ell h}$ by solving the following stochastic program
@@
$\begin{array}[]{cl}\min&\varphi(\widehat{\mathbb{P}}_{\ell h},\mathbb{Q}_{\ell h})\\ \mathrm{s.t.}&\mathbb{Q}_{\ell h}(\tilde{a}~{}\text{is classified by $\mathcal{C}_{\ell h}$ as 0})\geq 1-\gamma,~{}\mathbb{Q}_{\ell h}=\Delta_{\ell h}\#\widehat{\mathbb{P}}_{\ell h}.\end{array}$ (3)
@@
Problem (3) is easier to solve in specific circumstances. For example, when we impose that both ^$\widehat{\mathbb{P}}_{\ell h}$ and $\mathbb{Q}_{\ell h}$ are Gaussian and when we choose $\varphi$ as a moment-based divergence, then $\Delta_{\ell h}$ can be obtained by solving a convex optimization problem. In the next result, we use $\|\cdot\|_{F}$ as the Frobenius norm of a matrix, and $\Phi$ as the cumulative distribution function of a standard Gaussian distribution.
@@
Suppose that ^$\widehat{\mathbb{P}}_{\ell h}\sim\mathcal{N}(\widehat{\mu},\widehat{\Sigma})$ and $\mathbb{Q}_{\ell h}\sim\mathcal{N}(\mu,\Sigma)$ and $\varphi$ admits the form
@@
Proof of Theorem 1.
The logistic classifier $\mathcal{C}_{\ell h}$ output a prediction 0 if $\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h}<0$. If $\mathbb{Q}_{\ell h}$ is Gaussian $\mathcal{N}(\mu,\Sigma)$, then by [34, Theorem 10.4.1], the probability constraint of (3) can be written as
$\begin{array}[]{cl}\min&\|\mu-\widehat{\mu}\|_{2}^{2}+\|\Sigma^{\frac{1}{2}}-\widehat{\Sigma}^{\frac{1}{2}}\|_{F}^{2}\\ \mathrm{s.t.}&\vartheta_{\ell h}+\theta_{\ell h}^{\top}\mu+\Phi^{-1}(1-\gamma)t\leq 0,\quad\sqrt{\theta_{\ell h}^{\top}\Sigma\theta_{\ell h}}\leq t\\ &\mu\in\mathbb{R}^{d},~{}\Sigma\in\mathbb{S}_{+}^{d},~{}t\in\mathbb{R}_{+}.\end{array}$
Let $S\leftarrow\Sigma^{\frac{1}{2}}\in\mathbb{S}_{+}^{d}$, the constraint $\sqrt{\theta_{\ell h}^{\top}\Sigma\theta_{\ell h}}\leq t$ is equivalent to $\|S\theta_{\ell h}\|_{2}\leq t$, which leads to (4). Thus, the optimal pushforward $\Delta_{\ell h}$ should push ^$\widehat{\mathbb{P}}_{\ell h}\sim\mathcal{N}(\widehat{\mu},\widehat{\Sigma})$ to $\mathbb{Q}_{\ell h}\sim\mathcal{N}(\mu^{\star},(S^{\star})^{2})$. One can verify through simple linear algebraic calculations that the mapping $\Delta_{\ell h}(a_{\ell h})=G^{\star}_{\ell h}a_{\ell h}+g^{\star}_{\ell h}$ defined in the theorem statement is the desired mapping. This completes the proof. ∎
@@
Figure 2. Headwise intervention: at head $h$ of layer $\ell$, we learn a linear mapping $\Delta_{\ell h}$ that transforms the undesirable-predicted activations to desirable-predicted activations.
The effect of the headwise intervention $\Delta_{\ell h}$ is illustrated in Figure 2. The headwise classifier $\mathcal{C}_{\ell h}$ is represented by the red linear hyperplane $\vartheta_{\ell h}+\theta_{\ell h}^{\top}a=0$ on the activation space; the undesirable-predicted (label 1) region is towards the top left corner, while the desirable-predicted (label 0) region is towards the bottom right corner. The activations of the undesirable-predicted samples are represented as a Gaussian distribution with mean $(\widehat{\mu},\widehat{\Sigma})$, drawn as the red ellipsoid. The edit map $\Delta_{\ell h}$ pushes this distribution to another Gaussian distribution $\mathbb{Q}_{\ell h}$ drawn as the green ellipsoid. The distribution $\mathbb{Q}_{\ell h}$ has a coverage guarantee on the desirable-predicted region with probability at least $1-\gamma$. One can also verify that $\mathbb{Q}_{\ell h}$ has mean $\mu^{\star}$ and covariance matrix $(S^{\star})^{2}$. Problem (4) can be solved using semidefinite programming solvers such as COPT or Mosek.
@@
The moment information $\widehat{\mu}$ and $\widehat{\Sigma}$ can be estimated from the subset $\hat{\mathcal{D}}_{\ell h}^{+}$. One can intuitively expect a trade-off between the tolerance level $\gamma$ and the magnitude of the headwise mapping. If $\gamma$ is lower, the activations will be edited at a higher magnitude so that the edited activations will likely end up in the desirable-predicted region of the classifier $\mathcal{C}_{\ell h}$. In contrast, if $\gamma$ is higher, the activations will be edited with a smaller magnitude due to the lower stringent constraint to swap the predicted label.
@@
One can view the distribution $\mathbb{Q}_{\ell h}\sim(\mu^{\star},(S^{\star})^{2})$ as the counterfactual distribution of the undesirable-predicted activations with minimal perturbation. This distribution $\mathbb{Q}_{\ell h}$ is found by optimization, which is in stark contrast with the design of the counterfactual distribution in MiMic [37], in which the intervention is computed based on the activations of the desirable-predicted activations. As a comparison to ITI [24], we note that the headwise intervention of ITI does not depend on the value of the activations: ITI shifts the activations along the truthful directions for a stepsize multiplied by the standard deviation of activations along the intervention (truthful) direction. In contrast, our headwise intervention depends on the value $a_{\ell h}$, and one can verify that the magnitude of the proposed shift amounts to $\|(G^{\star}_{\ell h}-I)a_{\ell h}+g^{\star}_{\ell h}\|_{2}$. Moreover, ITI does not provide any (probabilistic) guarantee for the intervention, while the probabilistic guarantee is internalized in our method through the design of the map in equation (3).
@@
$\hat{a}_{\ell h}=\mathbbm{1}_{\mathcal{C}_{\ell h}(a_{\ell h})=1\text{ and }\mathcal{C}_{\ell}(a_{\ell})=1}\Delta_{\ell h}(a_{\ell h})\qquad\forall h=1,\ldots,H,$ (5)

### v2

# Distributional Surgery for Language Model Activations

@@ Abstract
Language models, while capable of generating remarkably coherent and seemingly accurate text, can occasionally produce undesirable content, including harmful or toxic outputs. In this paper, we present a new two-stage approach to detect and mitigate undesirable content generations by rectifying activations. First, we train an ensemble of layerwise classifiers to detect undesirable content using activations by minimizing a smooth surrogate of the risk-aware score. Then, for detected undesirable contents, we propose layerwise distributional steering policies that transform the attention heads. These policies are computed through principled semidefinite programming, which aims to minimally perturb the attention distribution while probabilistically guaranteeing the effectiveness of the editions. Empirical evaluations across multiple language models and datasets show that our method outperforms baselines in reducing the generation of undesirable output.
Distributional Surgery for Language Model Activations
Bao Nguyen CUHK nbnguyen@se.cuhk.edu.hk          Binh Nguyen VinUniversity binh.nt2@vinuni.edu.vn
Duy Nguyen UNC-Chapel Hill duykng@cs.unc.edu          Viet Anh Nguyen CUHK nguyen@se.cuhk.edu.hk
1 Introduction
Language models (LMs) have demonstrated a remarkable ability to understand and generate human-like documents Radford et al. (2019); Brown et al. (2020); Touvron et al. (2023a, b); Jiang et al. (2023); Dubey et al. (2024). However, inspection of their output often reveals undesirable generation, including inaccurate or toxic texts Ji et al. (2023); Rawte et al. (2023); Xu et al. (2024). Meanwhile, developing effective strategies to control the generation process of LMs remains a significant challenge Tonmoy et al. (2024).
@@ 1 Introduction
Researchers have proposed numerous methods for controllable text generation in language models Zhang et al. (2023); Li et al. (2024a). These approaches primarily include model editing and supervised fine-tuning. Both methods, however, require altering the model weights using a subset of text samples, which can result in unstable representations for other text instances Hase et al. (2024). Additionally, these approaches typically require substantial computational resources.
@@
To address these limitations, activation intervention has emerged as a promising alternative for controllable text generation Subramani et al. (2022); Hernandez et al. (2023); Li et al. (2024b). This approach involves altering the model activations responsible for undesirable output during inference. Previous research has identified interpretable directions within the activation space of language models that play a causal role during inference. Studies by Burns et al. (2022) and Moschella et al. (2023) suggest that these directions can be manipulated to adjust model behavior in a controlled manner. This body of work suggests that the internal representations of language models are structured in a way that enables fine-grained control over generated text.
Drawing inspiration from these findings, activation intervention frameworks operate on the premise that the information needed to guide the model toward generating a target sentence is already encoded within the model. The hidden information is extracted as latent vectors and then used to steer the generation toward producing desirable outputs. The preliminary success of these activation intervention methods motivates our approach to improving the quality and controllability of language model generation.
@@
Problem Statement. We consider a language model consisting of $L$ layers, each layer has $H$ heads, each head has dimension $d$. For example, for Llama-2, we have $L=32$, $H=32$, and $d=128$. The training dataset is denoted by $\mathcal{D}=(x_{i},y_{i}^{*})_{i=1,\ldots,N}$, the $i$-th text is denoted by $x_{i}$, and its ground truth label is $y_{i}^{*}\in\{0,1\}$, where the label 1 (positive) represents the undesirable text, and the label 0 (negative) represents the desirable text. Our goal is twofold: (i) detect an undesirable text, and (ii) modify an undesirable text into a desirable text.
@@
$a_{\ell+1,i}$ $=a^{\mathrm{mid}}_{\ell,i}+\mathrm{FFN}(a^{\mathrm{mid}}_{\ell,i}),$ (1)
$a^{\mathrm{mid}}_{\ell,i}$ $=a_{\ell,i}+\sum_{i=1}^{H}Q_{\ell h}\textrm{Att}(P_{\ell h}a_{\ell,i}).$
Here, $P_{\ell h}\in\mathbb{R}^{d\times dH}$ is the projection matrix that maps the output of each layer to the $d$ dimensional head space, $\mathrm{Att}$ is the attention operator Vaswani et al. (2017), $Q_{\ell h}\in\mathbb{R}^{dH\times d}$ is the pull-back matrix and $\mathrm{FFN}$ is the feed-forward layer. Each $a_{\ell,i}$ is a concatenation of headwise activations $a_{\ell h,i}$ for $h=1,\ldots,H$. Inspired by Li et al. (2024b), we aim to perform intervention at some selected $a_{\ell h,i}$, the activations for head $h$ of layer $\ell$, if we detect that the activation is from undesirable content.
Contributions. We contribute an activation intervention method to detect and rectify the undesirable generation of LM. We call our method RADIANT (Risk-Aware Distributional Intervention Policies for Language Models’ Activations). RADIANT comprises two components:
  1. 1.
@@ 1.1 Related Works
Controllable generation. Controllable text generation methods aim to alter the outputs of large language models in a desired way. One possible approach is model editing Wang et al. (2023a); Zhang et al. (2024), which involves modifying the parameters of a model to steer its outputs. For example, Meng et al. (2022) involves identifying specific middle-layer feedforward modules that correspond to factual knowledge and then altering these weights to correct or update the information encoded by the model. Other notable methods include fine-tuning techniques such as Supervised Fine-Tuning (SFT, Peng et al. 2023; Gunel et al. 2020) and Reinforcement Learning from Human Feedback (RLHF, Ouyang et al. 2022a; Griffith et al. 2013).
@@
Probing. Probing is a well-established framework to assess the interpretability of neural networks Alain and Bengio (2016); Belinkov (2022). Probing techniques have been applied to understand the internal representations of transformer architectures in language models such as BERT and GPT. For example, Burns et al. (2022) proposed an unsupervised probing method that optimizes consistency between positive and negative samples. Marks and Tegmark (2023) computes the mean difference between true and false statements and skews the decision boundary by the inverse of the covariance matrix of the activations.
@@
Activation interventions. Activation intervention at inference time is an emerging technique for controllable generation Turner et al. (2023); Li et al. (2024b); Singh et al. (2024); Yin et al. (2024). Unlike model editing or fine-tuning techniques, inference time intervention does not require altering the model parameters. Li et al. (2024b) proposed a headwise intervention method for eliciting truthful generated answers of a language model. They first train linear probes on each head of the language model, then shift the activations with the probe weight direction or the mean difference direction.
@@
There is a clear distinction between our method and ITI when choosing the location of the classifiers and, hence, the location of the interventions. The ITI method builds different headwise classifiers scattered at different layers, and it may suffer from distribution shifts: if an activation is intervened, this leads to shifts in the activation values at all subsequent layers in the network. Thus, classifiers trained at subsequent layers can degrade performance, and interventions at subsequent layers can also degrade. On the contrary, we build a layerwise classifier focusing on all heads in the same layer and do not suffer from the distributional shifts of the activations.
@@
The recent paper by Singh et al. (2024) is closely related to our work. The authors propose a heuristic intervention rule; then, using empirical estimations of the means and covariances of activations data’s distributions of desirable and undesirable text, they calculate a closed-form optimal transport plan between these two empirical distributions, assuming they are standard normal. However, this framework does not take into account the semantics of sentences. Another recent method, called LoFit (Localized Fine-Tuning on LLM Representations Yin et al. 2024), also identifies a specific subset of attention heads that are crucial for learning a particular task, but then performs fine-tuning on the intervention vectors at those chosen heads to enhance the model’s hidden representations. This results in additional training overhead.
2 Layerwise Risk-aware Probes
@@
[…]
The training process of $\mathcal{C}_{\ell h}$ must take into account two types of risk: (i) false-negative risk when an undesirable text is not detected, (ii) false-positive risk when a desirable text is classified as undesirable, and is subsequently edited and loses its original semantics. Therefore, a natural candidate for the loss function is a combination of the False Positive Rate (FPR) and the False Negative Rate (FNR). However, neither FPR nor FNR have smooth functions in optimizing variables. We, hence, resort to smooth surrogates of these two metrics that use the predicted probability of the classifier, similarly to Bénédict et al. (2022). In detail, we use
$\mathrm{FPR}(\theta_{\ell h},\vartheta_{\ell h})$
$=\frac{1}{N_{0}}\sum_{i=1}^{N}\mathrm{sigmoid}(\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h,i})\times(1-y^{*}_{i}),$
$\mathrm{FNR}(\theta_{\ell h},\vartheta_{\ell h})$
$=\frac{1}{N_{1}}\sum_{i=1}^{N}\left(1-\mathrm{sigmoid}(\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h,i})\right)\times y^{*}_{i}.$
@@
where $\tau\in[0,H]$ is a tunable threshold. When $\tau=\lfloor H/2\rfloor$, then $\mathcal{C}_{\ell}$ becomes the majority voting results of the individual (weak) classifiers $\mathcal{C}_{\ell h}$. We optimize the hyperparameter $\tau$ to reduce the False Negative Rate (FNR), with a secondary focus on the False Positive Rate (FPR) in cases where the FNR rates are equal. The rationale for this choice is that we believe undesirable content being labeled as desirable is more problematic than other instances.
@@
To conclude this step, we compute the classifier $\mathcal{C}_{\ell}$ for layer $\ell=1,\ldots,L$ by tuning the parameters $\alpha$. The layer whose classifier $\mathcal{C}_{\ell}$ delivers the highest quality (accuracy or any risk-aware metric) will be the optimal layer to construct the probe. This optimal layer, along with the collection of headwise classifiers, is the final output of this step.
3 Headwise Interventions with Probabilistic Guarantees
@@
We propose a distributional intervention to the activations of the samples predicted as undesirable by the layerwise classifier. In this section, we will focus on constructing a single headwise intervention, and in the next section, we will combine multiple headwise interventions into a layerwise intervention. A headwise intervention is a map $\Delta_{\ell h}:a_{\ell h}\mapsto\hat{a}_{\ell h}$ that needs to balance multiple criteria: (i) it should be easy to compute and deploy, (ii) it should be effective in converting the undesirable activations to the desirable regions, (iii) it should minimize the magnitude of the intervention to sustain the context of the input. Intuitively, we propose solving an optimization problem with the loss and constraints that fit all the criteria listed. The details are as follows.
@@
To promote (i), we employ a simple linear map $\Delta_{\ell h}(a_{\ell h})=G_{\ell h}a_{\ell h}+g_{\ell h}$ parametrized by a matrix $G_{\ell h}\in\mathbb{R}^{d\times d}$ and a vector $g_{\ell h}\in\mathbb{R}^{d}$. This linear map can also be regarded as a pushforward map that transforms the undesirable-predicted activations to become desirable-predicted activations. Let us now represent the undesirable-predicted activations as a $d$-dimensional random vector $\tilde{a}_{\ell h}$. Its distribution can be estimated using the training data after identifying the subset $\hat{\mathcal{D}}_{\ell h}^{+}$ of training samples that are predicted undesirable by $\mathcal{C}_{\ell h}$, that is, $\hat{\mathcal{D}}_{\ell h}^{+}\triangleq\{i:\mathcal{C}_{\ell h}(a_{\ell h,i})=1\}$. The activations of samples in $\hat{\mathcal{D}}_{\ell h}^{+}$ lead to an empirical distribution $\widehat{\mathbb{P}}_{\ell h}$. The linear map $\Delta_{\ell h}$ will pushforward the distribution $\widehat{\mathbb{P}}_{\ell h}$ to the new distribution $\mathbb{Q}_{\ell h}=\Delta_{\ell h}\#\widehat{\mathbb{P}}$.
@@
Using the pushforward distribution $\mathbb{Q}_{\ell h}$, we can impose criteria (ii) and (iii) above in an intuitive method. To promote (ii), we require that the activations distributed under $\mathbb{Q}_{\ell h}$ should be classified as desirable by $\mathcal{C}_{\ell h}$ with high probability. Finally, to promote (iii), we require that the distributions $\mathbb{Q}_{\ell h}$ and $\widehat{\mathbb{P}}_{\ell h}$ be not too far from each other. Let $\gamma\in(0,0.5)$ be a small tolerance parameter, and let $\varphi$ be a measure of dissimilarity between probability distributions, we propose to find $\Delta_{\ell h}$ by solving the following stochastic program
@@
$\begin{array}[]{cl}\min&\varphi(\widehat{\mathbb{P}}_{\ell h},\mathbb{Q}_{\ell h})\\ \mathrm{s.t.}&\mathbb{Q}_{\ell h}(\tilde{a}~\text{classified by $\mathcal{C}_{\ell h}$ as 0})\geq 1-\gamma,\\ &\mathbb{Q}_{\ell h}=\Delta_{\ell h}\#\widehat{\mathbb{P}}_{\ell h}.\end{array}$ (3)
@@
Problem (3) is easier to solve in specific circumstances. For example, when we impose that both $\widehat{\mathbb{P}}_{\ell h}$ and $\mathbb{Q}_{\ell h}$ are Gaussian and when we choose $\varphi$ as a moment-based divergence, then $\Delta_{\ell h}$ can be obtained by solving a convex optimization problem. In the next result, we use $\|\cdot\|_{F}$ as the Frobenius norm of a matrix, and $\Phi$ as the cumulative distribution function of a standard Gaussian distribution.
@@
[…]
@@
Figure 1: At head $h$ of layer $\ell$, we learn a headwise intervention, linear mapping $\Delta_{\ell h}$ to transform the undesirable-predicted activations to desirable-predicted activations.
[…]
@@
The moment information $\widehat{\mu}$ and $\widehat{\Sigma}$ can be estimated from the subset $\hat{\mathcal{D}}_{\ell h}^{+}$. One can intuitively expect a trade-off between the tolerance level $\gamma$ and the magnitude of the headwise mapping. If $\gamma$ is lower, the activations will be edited at a higher magnitude so that the edited activations will likely end up in the desirable-predicted region of the classifier $\mathcal{C}_{\ell h}$. In contrast, if $\gamma$ is higher, the activations will be edited with a smaller magnitude due to the less stringent constraint to swap the predicted label.
@@
One can view the distribution $\mathbb{Q}_{\ell h}\sim(\mu^{\star},(S^{\star})^{2})$ as the counterfactual distribution of the undesirable-predicted activations with minimal perturbation. This distribution $\mathbb{Q}_{\ell h}$ is found by optimization, which is in stark contrast with the design of the counterfactual distribution in MiMic Singh et al. (2024), in which the intervention is computed based on the activations of the desirable-predicted activations. As a comparison to ITI Li et al. (2024b), we note that the headwise intervention of ITI does not depend on the value of the activations: ITI shifts the activations along the truthful directions for a stepsize multiplied by the standard deviation of activations along the intervention (truthful) direction. In contrast, our headwise intervention depends on the value $a_{\ell h}$, and one can verify that the magnitude of the proposed shift amounts to $\|(G^{\star}_{\ell h}-I)a_{\ell h}+g^{\star}_{\ell h}\|_{2}$. Moreover, ITI does not provide any (probabilistic) guarantee for the intervention, while the probabilistic guarantee is internalized in our method through the design of the mapping in equation (3).
@@
$\hat{a}_{\ell h}=\begin{cases}\Delta_{\ell h}(a_{\ell h})&\text{if }\mathcal{C}_{\ell h}(a_{\ell h})\times\mathcal{C}_{\ell}(a_{\ell})=1,\\ a_{\ell h}&\text{otherwise,}\end{cases}$